# Distance-coupling as an Approach to Position and Formation Control

### v2

@@ Abstract
Studies of coordinated motion in autonomous vehicle groups have become a significant topic of interest in recent years. In this paper, we study the case when the agents are limited to distance measurements from predetermined reference points, or anchors. A novel approach, referred to as distance-coupling, is proposed for approximating agent positions from anchor distances. By taking the difference between squared distance functions, we cancel out the quadratic term and obtain a linear function of the position for the point of interest. We apply this method to the homing problem and prove Lyapunov stability with and without anchor placement error; defining bounds on the region of attraction when the anchors are linearly transformed from their desired positions. To further increase complexity, we show how the policy can be implemented on the formation control problem for a set of autonomous agents. In this context, we prove the existence of the set of equilibria for the distance-coupled formation controller and make preliminary claims on its convergence. The performance for both controllers are subsequently validated through simulation.
@@ I Introduction
Autonomously operated vehicles are commonly designed around high-cost measurement devices used for real-time localization. However, there exists instances, especially in low-cost systems, where the measurements available to the agents are limited to basic geometric information. We study the case when the information available is restricted to its Euclidean distance from $p$ fixed points, or anchors. Exploiting a linear, first-order model, we first address the homing problem. In this case, the primary goal is to move an agent from an arbitrary initial position to some predetermined equilibrium position using the available measurement data. Despite being based on nonlinear measurements, our controller is a simple linear function, and its stability can be proven in a straight-forward manner.
We also consider the problem of driving $m=p$ agents into some predetermined formation using only inter-agent distances. In this case, we show that the controller defined for the homing problem can be implemented with minimal adjustments. Subsequently, the set of equilibrium positions are proven through transformation invariance, and validated through simulation.
@@ I-A Literature Review
Studies of distance-based position functions have primarily been investigated through barycentric coordinates [1]. These methods represent the positions of anchors as vertices of a polygon and utilize its geometry to derive the agent’s position. While this method is robust in many applications, the point is usually limited to the bounds of a convex polyhedron and maintains ill-defined behaviors when this assumption is broken. To overcome this, researchers have implemented policy switching schemes to the formation control problem with success in 2-D [2] and 3-D [3] coordinate frames.
For distance-based formation control, the most popular approach uses gradient-descent over a function defined on the measurement graph [4, 5]. The graph can take many forms (directed/undirected, fully connected, etc.), but the connections are predetermined and static. Papers by Lin, Dimarogonas and Choi are also among the first to demonstrate the stability of assorted gradient-based formation control policies through standard analysis methods [6, 7, 8].
Another noteworthy example is that of [9] which is among the first written on the topic and discusses a catalog of analytical, methods which can be used to stabilize formations within small perturbations of their equilibrium. Lastly, the papers [10, 11] discuss the formation control of second-order systems, and [12] defines a policy in terms of the complex Laplacian. The review paper [13] lays a good foundation for many of the methods available. In addition to formation control in a distance-based context, it also discusses methods for the position and displacement-based measurement cases.
@@ I-B Paper Contributions
In this paper, we present a new method that algebraically cancels the nonlinear term from available distance functions. By taking the difference between squared-distance measurements, we show that we can accurately construct a linear approximation of the agent’s position in the environment. We refer to this method as distance-coupling.
The distance-coupled approach is novel in comparison to the literature as it does not have strict requirements on agent positions similar to those found in barycentric coordinate methods. Where preexisting methods primarily depend on the convexity of the anchor environment, our approach makes no such claims and allows the user to select anchor positions which are non-convex. However, we discuss degenerate collinearity cases and the minimum number of anchors required (Section III-A).
The new approach also enjoys better convergence expectations then the gradient-based alternative. For the homing problem, we prove that the linear controller exhibits global exponential convergence in the ideal case and a sufficiently large region of attraction when the anchors are linearly offset from their desired positions. For the formation control problem, we formally characterize the set of all equilibria through the controller’s transformation invariance, and confirm large regions of attraction empirically.
While we demonstrate our results in an $\mathbb{R}^{2}$ environment for simplicity, the methods discussed in Section III are generalized to $\mathbb{R}^{n}$.
@@ II System Definition
We start by defining the agent dynamics and ideal controller, along with the measurements available to the system.
@@ II-A System Dynamics
The agents are modeled as a linear, time-invariant single integrator of the form
$\dot{x}=u,$ (1)
where $x\in\mathbb{R}^{n}$ and $u\in\mathbb{R}^{n}$ are the state and the control input, respectively. As common in the literature, we assume that the environment is unbounded and without obstacles. The goal of the homing problem is to move an agent from some initial condition to an equilibrium position denoted by $x^{(\text{eq})}$.
For the case that there are $m$-agents operating in a single environment, we can write (1) in matrix form such that
$\dot{X}=U,$ (2)
where $X\in\mathbb{R}^{n\times m}$ and $U\in\mathbb{R}^{n\times m}$ are the state and control for the set of agents defined by
$X=\begin{bmatrix}x_{1}&\cdots&x_{m}\end{bmatrix}\text{, and }U=\begin{bmatrix}u_{1}&\cdots&u_{m}\end{bmatrix}.$ (3)
II-B Ideal Controller
Since the system is linear, we define a simple state feedback controller that brings the agent to the goal position $x^{(\text{eq})}\in\mathbb{R}^{n}$ as
$u(x)=C(x-x^{(\text{eq})}),$ (4)
where $C\in\mathbb{R}^{n\times n}$ describes the policy coefficients. By standard linear system theory, if $C$ has eigenvalues with negative real parts, then the system is globally exponentially stable and the state will converge to the desired equilibrium from any initial condition [14].
Unfortunately, we are limited to the measurements described in the next section and cannot define a direct controller from the state of the agent. That said, (4) will be referred to as the ideal control.
II-C Anchor and Anchor-distance Sets
Figure 1: Anchor set where $p=3$ so that $\mathcal{A}=\{a_{1},a_{2},a_{3}\}$ and $\mathcal{D}(x,A)=\{d(x,a_{1}),d(x,a_{2}),d(x,a_{3})\}$.
@@ II-C Anchor and Anchor-distance Sets
$\mathcal{A}=\{a_{i}\in\mathbb{R}^{n}:\forall i\leq p\},$ (5)
where $a_{i}$ is the position of the $i$-th anchor. We assume that the agent measures the distance from its position to each of the anchors. The anchor-distance set is thus defined as
@@
where
$d(x,a_{i})=\sqrt{(x-a_{i})^{\intercal}(x-a_{i})}.$ (7)
An example for when $p=3$ is shown in Figure 1. In the interest of conciseness, the individual distance functions from the point $x$ are referred to as $d_{i}(x)=d(x,a_{i})$, and the anchor-distance set is referred to as $\mathcal{D}(x)=\mathcal{D}(x,\mathcal{A})$.
@@ III Distance-coupled Position Control
In this section we show how (4) can be indirectly implemented as a function of the measurement set $\mathcal{D}(x)$ via distance-coupled functions. We then prove Lyapunov stability, characterize the region of attraction when the anchors are linearly transformed and show the behavior of the closed-loop system in simulation.
III-A Distance-coupled Position and Control
We start by squaring and expanding each element of the anchor-distance set (6) which results in
@@ III-A Distance-coupled Position and Control
The key step in our approach is to note that by subtracting the squared-distance functions for two anchors, the quadratic terms $x^{\intercal}x$ cancel out such that
@@
where $\mathcal{I}$ is the set of $(i,j)$ indices which make up every combination of the anchors included in (5). The vstack()($\cdot$)( ⋅ ) function simply converts a set of values into a vertically-oriented vector/matrix as appropriate. Using $|\mathcal{I}|$ to denote the cardinality of $\mathcal{I}$, we then have that $A\in\mathbb{R}^{|\mathcal{I}|\times n}$, and the vectors $b,h(x)\in\mathbb{R}^{|\mathcal{I}|}$.
Combining these with (10), the distance-coupled position of the agent can be computed by solving the linear system:
$Ax=h(x)-b$ (12)
$\Rightarrow A^{\intercal}Ax=A^{\intercal}(h(x)-b)$
[…]
where $K=(A^{\intercal}A)^{-1}A^{\intercal}$. Since the matrix $A^{\intercal}A$ is by definition positive definite, it is invertible if $A$ has full column rank with a subset of $n+1$ noncollinear anchor positions. While we do not include the proof here for conciseness, it is trivial and follows from standard linear algebra theory.
We refer to (12) as the distance-coupled position and can substitute it into (4) to obtain the corresponding distance-coupled controller:
@@
where $u(x)$ is solely dependent on the sets $\mathcal{A}$ and $\mathcal{D}(x)$.
@@ III-B Global Position Stability
We use Lyapunov’s method to prove stability of the distance-coupled controller (13) for an agent with model (1). Excluded from this paper for conciseness, a thorough definition of the Lyapunov stability criterion and the appropriate proofs can be found in [14]. We use $\mathbf{I}\in\mathbb{R}^{n\times n}$ to denote the identity matrix.
@@
The distance-coupled control policy (13) is globally exponentially stable for an agent with dynamics (1) if $C=\alpha\mathbf{I}$ given that $\alpha<0$.
@@
Define the Lyapunov candidate using the inner product between $x$ and $x^{(\text{eq})}$:
$V(x)=(x-x^{(\text{eq})})^{\intercal}(x-x^{(\text{eq})}).$ (14)
By definition, $V(x)=0$ at $x=x^{(\text{eq})}$ and is positive definite for all $x\neq x^{(\text{eq})}$, thus satisfying the first component to Lyapunov stability. Taking its derivative and substituting (13) for the control gives
$\dot{V}(x)$ $=2(x-x^{(\text{eq})})^{\intercal}\dot{x}=2(x-x^{(\text{eq})})^{\intercal}u(x)$ (15)
$=2(x-x^{(\text{eq})})^{\intercal}C(K(h(x)-b)-x^{(\text{eq})})$
Finally, using (12):
$\dot{V}(x)=2(x-x^{(\text{eq})})^{\intercal}C(x-x^{(\text{eq})}).$ (16)
Implying that $\dot{V}=0$ when $x=x^{(\text{eq})}$ and $\dot{V}<0$ given that $C=\alpha\mathbf{I}\prec 0$ with $\alpha<0$. Thus making the controller globally exponentially stable via Lyapunov stability [14].
Note that the proof can be extended to the case where $C$ is any stable matrix. However, we use the restriction $C=\alpha\mathbf{I}$ for the sake of analysis in the next section.
@@ III-C Stability Under Transformations of the Anchor Set
Let us now study the case when the set $\mathcal{A}$ is transformed by some orthogonal rotation, $R\in\mathbb{R}^{n\times n}$, and offset $r\in\mathbb{R}^{n}$, but the distance-coupled controller components (11) remain constant. That is, $\tilde{a}_{i}\in\mathbb{A}$ are the coordinates which were originally used to derive (11), where $\mathbb{A}$ is referred to as the anchor frame. The true positions are denoted $a_{i}\in\mathbb{W}$, where $\mathbb{W}$ is the world frame.
For a given anchor we can thus define the transformation
$a_{i}=R\tilde{a}_{i}+r:\tilde{a}_{i}\in\mathcal{A}.$ (17)
We differentiate between the two frames as the transformation of the set $\mathcal{A}$ creates error in the agent position defined by (12). That is,
$\tilde{x}$ $=K(h(x)-b)=R^{\intercal}(x-r),$ (18)
where $x\in\mathbb{W}$ represents the world frame position of the agent and $\tilde{x}\in\mathbb{A}$ is its approximation in the anchor frame. The distance-coupled control function similarly becomes
$\tilde{u}(x)$ $=C(K(h(x)-b)-x^{(\text{eq})})$ (19)
$=C(\tilde{x}-x^{(\text{eq})})=C(R^{\intercal}(x-r)-x^{(\text{eq})})$
$=R^{\intercal}C(x-(Rx^{(\text{eq})}+r))$
Likewise to Proposition 1, we can prove stability in terms of a range of acceptable values for the rotation and offset of the anchors. In order to make concrete claims on the stability of a system in $\mathbb{R}^{2}$, we identify the range of acceptable $\theta$ which make $R$ positive definite. This finding will be subsequently used in Proposition 2.
@@
A 2-D rotation defined by $R(\theta)\in\mathbb{R}^{2\times 2}:\theta\in\mathbb{R}$ is positive definite for the range $\theta\in(2\pi k-\tfrac{\pi}{2},2\pi k+\tfrac{\pi}{2})$ $\forall k\in\mathbb{Z}$ where $\mathbb{Z}$ is the set of integers.
@@
Define the matrix $R$ in terms of Rodrigues’ formula for 2-D rotations:
$R(\theta)=\begin{bmatrix}\cos(\theta)&-\sin(\theta)\\ \sin(\theta)&\cos(\theta)\end{bmatrix}$ (20)
where $\theta$ is some scalar angle. For some arbitrary point, $x$, we can say $R\succ 0$ when $x^{\intercal}Rx>0$ for all $x\neq 0$. Expanding this condition we get that
$x^{\intercal}Rx$ $=[x_{1}\ x_{2}]\begin{bmatrix}\cos(\theta)&-\sin(\theta)\\ \sin(\theta)&\cos(\theta)\end{bmatrix}\begin{bmatrix}x_{1}\\ x_{2}\end{bmatrix}>0$ (21)
$=(x_{1}^{2}+x_{2}^{2})\cos(\theta)>0,$
letting us conclude that $R(\theta)$ is positive definite so long as $\cos(\theta)>0$. This is true for the ranges
$\theta\in(2\pi k-\tfrac{\pi}{2},2\pi k+\tfrac{\pi}{2})\ \forall k\in\mathbb{Z},$ (22)
where $\mathbb{Z}$ is the set of integers.
By this, the stability candidate developed in Proposition 1 can be reformulated for the rotated anchor case and used to identify the region of attraction for (19).
@@
[…]
@@
Start by assuming that the new equilibrium, $\tilde{x}^{(\text{eq})}\in\mathbb{R}^{2}$, takes an equivalent transformation as that defined over the set $\mathcal{A}$, making the new Lyapunov candidate:
@@
$=(x-(Rx^{(\text{eq})}+r))^{\intercal}(x-(Rx^{(\text{eq})}+r)).$
We thus have $V(x)=0:x=\tilde{x}^{(\text{eq})}=Rx^{(\text{eq})}+r$ and $V(x)>0$ otherwise. Taking the derivative we get
$\dot{V}(x)$ $=2(x-(Rx^{(\text{eq})}+r))^{\intercal}\tilde{u}(x)$ (24)
$=2(x-(Rx^{(\text{eq})}+r))^{\intercal}C(K(h(x)-b)-x^{(\text{eq})}),$
where $\tilde{u}(x)$ is the controller defined in (19). Thus giving
$\dot{V}(x)$ $=2(x-(Rx^{(\text{eq})}+r))^{\intercal}C(R^{\intercal}(x-r)-x^{(\text{eq})})$ (25)
$=2(x-(Rx^{(\text{eq})}+r))^{\intercal}R^{\intercal}C(x-(Rx^{(\text{eq})}+r)).$
By inspection, we can conclude that $\dot{V}(x)=0:x=\tilde{x}^{(\text{eq})}=Rx^{(\text{eq})}+r$. Moreover, $\dot{V}(x)$ is negative definite so long as $R^{\intercal}C\prec 0$. If we assume $C\prec 0$ by Proposition 1, then, $R^{\intercal}C\prec 0$ if $R$ is defined as in Lemma 1.
Remark:
While Proposition 2 is limited to $\mathbb{R}^{2}$ for simplicity, we identify that, given $C\prec 0$, the final claim holds for any system in $\mathbb{R}^{n}$ given that $R$ is a positive definite orthogonal rotation in $\mathbb{R}^{n}$.

### v3

# references.bib

@@ Abstract
In this letter, we study the case of autonomous agents which are required to move to some new position based solely on the distance measured from predetermined reference points, or anchors. A novel approach, referred to as distance-coupling, is proposed for calculating the agent’s position exclusively from differences between squared distance measurements. The key insight in our approach is that, in doing so, we cancel out the measurement’s quadratic term and obtain a function of position which is linear. We apply this method to the homing problem and prove Lyapunov stability with and without anchor placement error; identifying bounds on the region of attraction when the anchors are linearly transformed from their desired positions. As an application of the method, we show how the policy can be implemented for distributed formation control on a set of autonomous agents, proving the existence of the set of equilibria.
@@ I Introduction
Autonomously operated vehicles are commonly designed around high-cost measurement devices used for real-time localization. However, there exists instances, especially in low-cost systems, when the measurements available to the agents are limited to basic geometric information. We study the case where the agent’s information is restricted to the Euclidean distance from $p$ fixed points, or anchors. To demonstrate, we control a single agent with linear, first-order dynamics moving from an arbitrary initial position to an equilibrium position given by target reference distances. Despite being based on nonlinear measurements, our controller is a simple linear function, and its stability can be proven in a straightforward manner. We also consider the problem of driving $m=p$ agents into some predetermined formation using only inter-agent distances. We show that the controller defined for the homing problem can be implemented with minimal adjustments. The set of equilibrium positions are then proven and validated empirically.
@@ I-A Literature Review
Studies of distance-based position functions have primarily been investigated through barycentric coordinates [floater_generalized_2015]. These methods represent the positions of anchors as vertices of a polygon and utilize its geometry to derive the agent’s position. While this method is robust in many applications, the point is usually limited to the bounds of a convex polyhedron and maintains ill-defined behaviors when this assumption is broken. To overcome this, researchers have implemented policy switching schemes to the formation control problem with success in 2-D coordinate frames [fathian_distributed_2019].
For the formation control problem, the most popular approach implements gradient-descent with an objective function defined on the measurement graph [dimarogonas_connection_2008, dimarogonas_leaderfollower_2009]. The graph can take many forms (directed/undirected, fully connected, etc.), but the communication connections are fixed. Papers by Dimarogonas, and Choi are also among the first to demonstrate the stability of assorted gradient-based formation control policies through standard analysis methods [dimarogonas_stability_2008, choi_distance-based_2021].
The papers [ren_distributed_2007, chen_formation_2022] discuss the formation control of second-order systems, and [lin_distributed_2014] defines a policy in terms of the complex Laplacian. The review paper [oh_survey_2015] lays a good foundation for many of the methods available. In addition to formation control in a distance-based context, it also discusses methods for the position and displacement-based measurement cases. It discusses that, many instances the feedback laws require bearings or relative positions, so they are not purely distance-based.
@@ I-B Paper Contributions
In this letter, we present a new localization method which we refer to as distance-coupling. By taking the difference between squared-distance measurements, we show that we can accurately construct a linear approximation of the agent’s position in an environment given we know the positions of the reference points. Where preexisting methods depend on the convexity of the anchor environment and either bearings, or relative position information, our approach makes no such claims and allows the user to select anchor positions which are non-convex. However, we discuss degenerate collinearity cases and the minimum number of anchors required.
For the homing problem, we prove that the linear controller exhibits global exponential convergence in the ideal case and a large region of attraction when the anchors deviate from their expected positions. For the formation control problem, we formally characterize the set of all equilibria through the controller’s transformation invariance, and confirm large regions of attraction empirically. Finally, we demonstrate our results in an $\mathbb{R}^{2}$ environment; with methods trivially generalized to any dimension.
@@ II System Definition
We start by defining the agent dynamics, and the measurements available to the system at each point in time.
@@ II-A System Dynamics
The agents are modeled as a linear, time-invariant single integrator of the form $\dot{x}=u,$ where $x\in\mathbb{R}^{n}$ and $u\in\mathbb{R}^{n}$ are the state and the control input, respectively. As is common in the literature, we assume that the environment is unbounded, and without obstacles. We will represent an agent’s intended equilibrium position as $x^{(\text{eq})}$.
For the case where there are $m$ agents operating in a single environment, we can rewrite $\dot{x}$ in matrix form such that $\dot{X}=U$, where $X\in\mathbb{R}^{n\times m}$ and $U\in\mathbb{R}^{n\times m}$ are the state and control for the set of agents defined by $X=[x_{1}\cdots x_{m}]$, and $U=[u_{1}\cdots u_{m}]$.
II-B Anchor and Anchor-distance Sets
Figure 1: Anchor set with $p=3$ such that $\mathcal{A}=\{a_{1},a_{2},a_{3}\}$ and $\mathcal{D}(x,A)=\{d(x,a_{1}),d(x,a_{2}),d(x,a_{3})\}$.
@@ II-B Anchor and Anchor-distance Sets
$\mathcal{A}=\{a_{i}\in\mathbb{R}^{n}\ \forall i\leq p\},$ (1)
where $a_{i}$ is the position of the $i$-th anchor. We also assume that the agent measures the distance from its position to each of the anchors. The anchor-distance set is thus defined as
@@
where $d(x,a_{i})=\sqrt{(x-a_{i})^{\intercal}(x-a_{i})}.$
An example for when $p=3$ is shown in Figure 1. In future sections, we will also refer to the anchor distances as $d_{i}(x)=d(x,a_{i})$ and the set $\mathcal{D}(x)=\mathcal{D}(x,\mathcal{A})$. Again, we stress that the only measurements available to an agent is the set $\mathcal{D}(x)$, and their (possibly inexact) location with respect to the global environment.
@@ III Distance-coupled Position Control
In this section we show how a linear controller can be indirectly implemented as a function of the (nonlinear) measurement set $\mathcal{D}(x)$ via distance-coupled functions. We then prove Lyapunov stability, characterize the region of attraction when the anchors are linearly transformed, and show the behavior of the closed-loop system in simulation.
III-A Distance-coupled Position Functions
We start by squaring and expanding each element of the anchor-distance set (2):
@@ III-A Distance-coupled Position Functions
The key step in our approach is to note that by subtracting the squared-distance functions for two anchors, the nonlinear term $x^{\intercal}x$ cancels out such that
@@
Note that $a_{i}$, and $a_{j}$ are expected to be known, as opposed to $d_{i}$ and $d_{j}$ which are measured.
@@
where $\mathcal{I}$ is the set of $(i,j)$ indices which make up every combination of the anchors included in (1). The vstack()($\cdot$)( ⋅ ) function concatenates a set of values into a vertically-oriented vector/matrix as appropriate. Using $|\mathcal{I}|$ to denote the cardinality of $\mathcal{I}$ we have the matrix $A\in\mathbb{R}^{|\mathcal{I}|\times n}$, and the vectors $b,h(x)\in\mathbb{R}^{|\mathcal{I}|}$.
Combining (5) and (6), the distance-coupled position of the agent can be computed by solving the linear system:
$Ax=h(x)-b\Rightarrow A^{\intercal}Ax=A^{\intercal}(h(x)-b)$ (7)
$\Rightarrow x=K(h(x)-b),$
where $K=(A^{\intercal}A)^{-1}A^{\intercal}$. Since the matrix $A^{\intercal}A$ is always positive definite, it is invertible if $A$ has full column rank. This condition is satisfied if there exists a subset of at least $n+1$ noncollinear anchor positions.
We refer to (7) as the distance-coupled position. The corresponding distance-coupled controller is defined as
@@
where $C$ is a stable controller gain matrix, and $u(x)$ is solely dependent on the sets $\mathcal{A}$ (known) and $\mathcal{D}(x)$ (measured).
@@ III-B Global Position Stability
We use Lyapunov’s direct method to prove stability of the distance-coupled controller (8) for an agent with dynamics defined by the single integrator model [khalil_nonlinear_2001].
@@
The distance-coupled control policy (8) is globally exponentially stable for an agent with dynamics $\dot{x}=u$ if $C=\alpha\mathbf{I}$ with $\alpha<0$, and the identity $\mathbf{I}$.
@@
Define $V(x)=(x-x^{(\text{eq})})^{\intercal}(x-x^{(\text{eq})})$ as the positive definite Lyapunov candidate. Then, using $\dot{x}=u$ and controller (8) with $C=\alpha\mathbf{I}$ it is trivial to show that $\dot{V}=2\alpha(x-x^{(\text{eq})})^{\intercal}(x-x^{(\text{eq})})$. Implying that $\dot{V}=0$ when $x=x^{(\text{eq})}$ and otherwise $\dot{V}<0$ given that $\alpha<0$. Therefore, for all $\alpha<0$ the controller is globally exponentially stable via Lyapunov stability [khalil_nonlinear_2001].
Remark:
Note that Proposition 1 can be extended to the case where $C$ is any stable matrix. However, we use the restriction $C=\alpha\mathbf{I}$ for the sake of analysis in the next section.
@@ III-C Stability Under Transformations of the Anchor Set
We define the agent frame as $\mathbb{A}$, and the world frame $\mathbb{W}$. The controller (8) uses anchor positions $\tilde{a}_{i}\in\mathbb{A}$, but the measurements depend on the position $a_{i}\in\mathbb{W}$. The two sets are related by an unknown orthogonal rotation $R$ and offset $r$ such that $a_{i}=R\tilde{a}_{i}+r$. We identify that the transformation of the set $\mathcal{A}$ creates error in the agent position defined by (7), that is, $\tilde{x}=K(h(x)-b)=R^{\intercal}(x-r)$, where $x\in\mathbb{W}$ represents the world frame position of the agent and $\tilde{x}\in\mathbb{A}$ is its approximation in the anchor frame. Similarly, the controller (8) becomes
$\tilde{u}(x)$ $=C(K(h(x)-b)-x^{(\text{eq})}),$ (9)
$=C(\tilde{x}-x^{(\text{eq})})=C(R^{\intercal}(x-r)-x^{(\text{eq})}),$
$=R^{\intercal}C(x-(Rx^{(\text{eq})}+r)).$
In order to make stability claims similar to those discussed in Proposition 1, we will limit our system to $\mathbb{R}^{2}$, and identify the range of rotations which make $R\succ 0$.
@@
An orthogonal rotation $R\in\mathbb{R}^{2\times 2}$ defined by $\theta\in\mathbb{R}$ is positive definite in the ranges $\theta\in(2\pi k-\tfrac{\pi}{2},2\pi k+\tfrac{\pi}{2})$, where $k\in\mathbb{Z}$ and $\mathbb{Z}$ is the set of integers.
@@
Define the orthogonal rotation $R$ in terms of Rodrigues’ formula for 2-D rotations where the angle $\theta\in\mathbb{R}$. For some arbitrary point, $x$, we can say $R\succ 0$ when $x^{\intercal}Rx>0$ for all $x\neq 0$. After expanding, we get that $(x_{1}^{2}+x_{2}^{2})\cos(\theta)>0$; letting us conclude that $R$ is positive definite so long as $\cos(\theta)>0$. This is true for the ranges
$\theta\in(2\pi k-\tfrac{\pi}{2},2\pi k+\tfrac{\pi}{2})\text{ for all }k\in\mathbb{Z},$ (10)
where $\mathbb{Z}$ is the set of all integers.
Using Lemma 1, the stability candidate developed in Proposition 1 can be reformulated for the rotated anchor case and used to identify the region of attraction for (9).
@@
For a 2-D system where the set $\mathcal{A}$ has been transformed by an orthogonal rotation, $R$, and translation, $r$, the transformed distance-coupled controller (9) describes an exponentially stable policy with equilibrium $x=Rx^{(\text{eq})}+r$ for any $r$ and any $R$ such that $R^{\intercal}C\prec 0$ (where $C$ is defined in Proposition 1 and $R$ in Lemma 1).
@@
Define the transformed equilibrium $\tilde{x}^{(\text{eq})}\in\mathbb{R}^{2}$ so that $\tilde{x}^{(\text{eq})}=Rx^{(\text{eq})}+r$. Placing this into our original Lyapunov candidate we have,
@@
$=(x-(Rx^{(\text{eq})}+r))^{\intercal}(x-(Rx^{(\text{eq})}+r)),$
where $V=0$ when $x=Rx^{(\text{eq})}+r$ and $V>0$ otherwise. Taking the derivative of $V$ using (9) we get that
$\dot{V}(x)=2(x-(Rx^{(\text{eq})}+r))^{\intercal}R^{\intercal}C(x-(Rx^{(\text{eq})}+r)).$ (12)
By inspection, we can conclude that $\dot{V}(x)=0$ when $x=Rx^{(\text{eq})}+r$, and that $\dot{V}\prec 0$ so long as $R^{\intercal}C\prec 0$. If we assume $C\prec 0$ by Proposition 1, and $R$ is selected by the bounds defined in Lemma 1, then $R^{\intercal}C\prec 0$ if.